OpenStudy (anonymous):

Need help.. Dawn baked cookies & gave 3/4 of them to Penny. Penny gave back a dozen. Dawn ended up with 18 cookies. How many did she originally make?

OpenStudy (mathstudent55):

Dawn baked x cookies. Using x, how many did Dawn give to Penny?

OpenStudy (anonymous):

3/4

OpenStudy (welshfella):

No its (3/4)x cookies that he gave to penny

OpenStudy (welshfella):

No its 3/4 of x cookies that he gave to penny written as (3/4) x

OpenStudy (welshfella):

after giving (3/4)x to penny dawn , in terms of x, had how many copies?

OpenStudy (welshfella):

?? - those messages sure got mixed up!!!

OpenStudy (anonymous):

Yeahhh

OpenStudy (welshfella):

how many do you think dawn had left If you give away 3/4 of some quantity how much have you got left yourself?

OpenStudy (anonymous):

Dawn has 18 cookies left. To be honest I don't know the answer, 3/4 is .75

OpenStudy (welshfella):

you need to create an equation to solve for x If dawn originally had x cookies and she gave away 3/4 of them - that is (3/4)x , she would have x - (3/4)x = (1/4) x left. Do you get that?

OpenStudy (anonymous):

Not really

OpenStudy (welshfella):

(3/4)x simply means 3/4 of x what left after taking (3/4) x away from x is 1/4 of x , (1/4)x

OpenStudy (welshfella):

- its a kind of shorthand

OpenStudy (anonymous):

1/4 is .25

OpenStudy (welshfella):

if for instance x was 16 then 3/4 of it would be (3/4) 16 = 12 and (1/4) 16 = 4 would be left

OpenStudy (welshfella):

if you prefer you can use decimal x - 0.75x = 0.25x

OpenStudy (welshfella):

oh wel - I tried...

OpenStudy (mathstudent55):

Ok, here is the problem from the beginning: Dawn baked an unknown (to us) number of cookies. We will use the variable x to represent that unknown number. Then Dawn gave 3/4 of the number of cookies to Penny. Since we are letting x represent the number of cookies Dawn baked, that means Dawn gave 3/4 of x to Penny. 3/4 of x is the same as (3/4)x. After Dawn gave away 3/4 of the cookies she baked, she was left with 1/4 of the cookies she baked. We still don;t know how many cookies Dawn baked, so after giving the cookies to Penny, Dawn kept 1/4 of x cookies, or (1/4)x. Then Penny gave 12 cookies back to Dawn. Since Dawn still had (1/4)x cookies, now Dawn has (1/4)x + 12 cookies. We are told that Dawn now has 18 cookies. That means that (1/4)x + 12 = 18 We need to solve that equation for x to find how many cookies Dawn baked. (1/4)x + 12 = 18 Subtract 12 from both sides: (1/4)x = 6 Multiply both sides by 4: x = 24 Dawn originally baked 24 cookies. Now let's check our answer. Dawn baked 24 cookies. 1/4 of 24 is 6. 3/4 of 24 is 18. Dawn gave 3/4 of the 24 cookies to Penny. She gave 18 cookies away. Dawn kept 6 cookies. Then Penny gave 12 cookies back to Dawn. Now Dawn has 6 + 12 = 18 cookies, which is what the problem states, so our answer is correct.